# Fundamental examples of Reeb spaces of smooth functions defined from two graphs of smooth functions with same asymptotic behaviors

### v2

@@ Abstract.
Smooth, real analytic, or real algebraic (real polynomial) functions and maps. Reeb spaces. Graphs. Digraphs. Reeb graphs
@@
These tools or objects have been well-known as fundamental and important tools or objects in theory of Morse functions and more general differentiable functions with nice singularity ([14]). $R_{c}$ is, for a function $c$ which is not so wild mathematically, has the structure of a graph. For the Morse function case and Morse-Bott function case (on compact manifolds), see [2, 11] and for more general cases, see [1, 15, 16], for example. Especially, [1, Theorem 7.5], [15, Theorem 3.1], and [16, Theorems 2.1 and 2.8 and ”2 Reeb space and its graph structure”] are important as theory on topologies and graph structures of these spaces.
Hereafter, let ${\mathbb{R}}^{k}$ denote the $k$-dimensional Euclidean (real affine) space with $\mathbb{R}:={\mathbb{R}}^{1}$. Let ${\pi}_{k_{1},k_{2}}:{\mathbb{R}}^{k_{1}}\rightarrow{\mathbb{R}}^{k_{2}}$ denote the canonical projection mapping $x=(x_{1},x_{2})\in{\mathbb{R}}^{k_{1}-k_{2}}\times{\mathbb{R}}^{k_{2}}$ to $x_{1}$ with $k_{1}>k_{2}\geq 1$. A graph means a $1$-dimensional connected CW complex such that the closure of each $1$-cell (edge) is homeomorphic to the $1$-dimensional disk $D^{1}:=\{x\in\mathbb{R}\mid x^{2}\leq 1\}$: 0-cell is a vertex. As a wider class, a graph with ends means a $1$-dimensional connected CW complex such that the closure of each $1$-cell (edge) is homeomorphic to the $1$-dimensional disk $D^{1}:=\{x\in\mathbb{R}\mid x^{2}\leq 1\}$ or a ray $\{x\geq 0\mid x\in\mathbb{R}\}$: 0-cell is a vertex as above.
In the case of the smooth manifold $X$ with no boundary, $R_{c}$ is, in a certain nice case, a graph (with ends) whose vertex set consists of points for all critical contours of $c$. This is the so-called Reeb graph of $c$ and in most studies including presented studies above, this definition is adopted. We can see the Reeb graph as a digraph (with ends) by giving orientation on each edge by the following rule: an edge $e$ departing from a vertex $v_{e}$ is an oriented edge departing from (resp. entering) $v_{e}$.
@@
Recently the author has started to consider infinite graphs (with ends) as Reeb spaces and obtained several explicit cases. They are also constructed as, fully or densely, real algebraic or real analytic and based on the paper [5] and the preprints [7, 8]. Related preprints are [9, 10]. These preprints are important in our present paper, where we ignore related non-trivial knowledge and discuss related arguments in self-contained ways. The present paper is also a kind of our new remark on examples of [9, 10], especially [9, Theorem 1]. There real algebraic or real analytic functions or maps are presented and answers to Problem 1 are presented as the restrictions of the canonical projections ${\pi}_{k_{1},k_{2}}$ to the zero sets of real analytic or algebraic maps on the real affine spaces. We respect these examples and present answers from a kind of essentially generalized viewpoints.
In the next section, we prove our new propositions and theorems there. This is for a generalization of the case presented in [9, Theorem 2] and in Example 1, this is explained again. Explicitly, we consider a natural smooth map onto the region surrounded by the graphs of two smooth real-valued functions $c_{1}$ and $c_{2}$ goes to a same value at each infinity in ${\mathbb{R}}^{2}$, obtained via Theorem 1, and its composition with the canonical projection ${\pi}_{2,1}$.
@@
A proper map $c:X\rightarrow Y$ between topological spaces $X$ and $Y$ is a map whose preimage $c^{-1}(K)$ is compact for any each subset $K\subset X$. A non-proper map is one which is not proper.
@@
This means that the ranks of the differential of the function are always $1$ on the zero set. The zero set is an $m$-dimensional smooth submaifold with no boundary, We can prove the remaining easily by the situation. ∎
@@
    Suppose that the function $c_{i}:\mathbb{R}\rightarrow\mathbb{R}$ ($i=1,2$) is a smooth function the intersection of $\mathbb{R}-(\{q_{1}\}\bigcup\{q_{2}\})$ and whose critical value set $c_{i}(S(c_{i}))$ is a discrete set and at least one function of $c_{1}$ and $c_{2}$ has at least one critical point. Suppose also the following.
@@
      As $t\in\mathbb{R}$ goes to -$\infty$, then both the values $c_{1}(t)$ and $c_{2}(t)$ converge to $q_{1}\in\mathbb{R}$.
@@
      As $t\in\mathbb{R}$ goes to $\infty$, then both the values $c_{1}(t)$ and $c_{2}(t)$ converge to $q_{2}\in\mathbb{R}$.
@@
      For $i=1,2$, each connected component $K_{c_{1},c_{2},q_{i},j}$ of the set ${{\pi}_{m+1,1}}^{-1}(q_{i})\bigcap X_{m,c_{1},c_{2}}$ has a small compact and connected neighborhood $N_{c_{1},c_{2},q_{i},j}$ consisting of contours of the the restriction ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ and the complementary set of $K_{c_{1},c_{2},q_{i},j}$ in the connected component $N_{c_{1},c_{2},q_{i},j}$ consists of finitely many connected sets and contains no critical contour of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$. In addition, these neighborhoods $N_{c_{1},c_{2},q_{i},j}$ can be chosen to be mutually disjoint.
@@
We prove the statement (2). First, the critical sets of the functions $c_{1}$ and $c_{2}$ are discrete by the assumption. From the conditions (2a, 2b) and the structures of the maps and the manifolds, the restriction of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ to the preimage of $\mathbb{R}-(\{q_{1}\}\bigcup\{q_{2}\})$ is proper. The critical set of the restriction of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ to the preimage of each closed subset in $\mathbb{R}-(\{q_{1}\}\bigcup\{q_{2}\})$ is finite, and by virtue of [15, 16], its Reeb space is regarded to be homeomorphic to a $1$-dimensional manifold or a graph which is not homeomorphic to a manifold. In addition, if at least one critical point of the function exists, then the Reeb space of the restriction of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ to the preimage of $\mathbb{R}-(\{q_{1}\}\bigcup\{q_{2}\})$ is regarded to be the Reeb graph. For [15, 16], related to this argument, mainly, consult [15, Theorem 3.1] and [16, Theorems 2.1 and 2.8 and ”2 Reeb space and its graph structure”]. Related to this theory, see also [1]. We consider the condition (2c). According to this, around each point in ${\bar{{\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}}}^{-1}(\{q_{1}\}\bigcup\{q_{2}\})$, the Reeb space is homeomorphic to a $1$-dimensional manifold or a finite graph which is not homeomorphic to a manifold and regarded to be the Reeb graph (if there exists a critical point of the function ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ there), locally. This also comes from the general theory above.
@@
  • There exists an interval $\{t\mid a_{0}\leq t\leq b_{0}\}$ and $c$ has the maximum $c_{\rm M}$ there. Furthermore, there exist a sequence $\{a_{j}\}_{j=1}^{\infty}$ such that $a_{j}<a_{j+1}$ for each $j\geq 0$ and that as $j$ goes to $\infty$, $a_{j}$ goes to $\infty$ and another sequence $\{b_{j}\}_{j=1}^{\infty}$ such that $b_{j+1}<b_{j}$ for each $j\geq 0$ and that as $j$ goes to $\infty$, $b_{j}$ goes to -$\infty$ enjoying the following.
@@
    The case $c_{1}(x):=\frac{1}{2(x^{2}+1)}{\sin}^{2}x$ and $c_{2}(x):=\frac{1}{x^{2}+1}$ is an explicit case of Proposition 1 with the critical set of $c_{1}$ being unbounded, $c_{2}$ being (almost) unimodal, and $(q_{1},q_{2})=(0,0)$. This is also [9, Theorem 1].
@@
    The case $c_{1}(x):=-\frac{1}{2(x^{2}+1)}{\sin}^{2}x$ and $c_{2}(x):=\frac{1}{x^{2}+1}$ seems to be an explicit case of Proposition 1 with the critical set of $c_{1}$ being unbounded, $c_{2}$ being almost unimodal, and $(q_{1},q_{2})=(0,0)$. However, the Reeb space is not homeomorphic to any graph with ends. In fact, the condition (2c) from Theorem 2 is dropped and the set ${{\pi}_{2,1}}^{-1}(q_{i})\bigcap\overline{D_{c_{1},c_{2}}}$ is connected and non-compact, and incident to infinitely many $1$-cells: the Reeb space has the structure of a $1$-dimensional cell complex which is not a CW complex.
In Proposition 1, if the critical set of $c_{2}$ is bounded below and $c_{2}$ is almost-unimodal, then the restriction of $c_{2}$ to some interval $\{t\leq b_{\rm m}\}\subset\mathbb{R}$ is strongly increasing (for distinct numbers $t_{1}<t_{2}$ in the interval $c_{2}(t_{1})<c_{2}(t_{2})$), and if the critical set of $c_{2}$ is bounded above and $c_{2}$ is almost-unimodal, then the restriction of $c_{2}$ to some interval $\{t\geq b_{\rm M}\}\subset\mathbb{R}$ is strongly decreasing (for distinct numbers $t_{1}<t_{2}$ in the interval $c_{2}(t_{1})>c_{2}(t_{2})$). There if the critical set of $c_{1}$ is bounded below and $c_{1}$ is almost-unimodal, then the restriction of $c_{1}$ to some interval $\{t\leq b_{\rm m}\}\subset\mathbb{R}$ is strongly increasing (for distinct numbers $t_{1}<t_{2}$ in the interval $c_{1}(t_{1})<-c_{1}(t_{2})$), and if the critical set of $c_{1}$ is bounded above and $c_{1}$ is almost-unimodal, then the restriction of $c_{1}$ to some interval $\{t\geq b_{\rm M}\}\subset\mathbb{R}$ is strongly decreasing (for distinct numbers $t_{1}<t_{2}$ in the interval $c_{1}(t_{1})>-c_{1}(t_{2})$).
@@
Let $c_{1}$ and $c_{2}$ be functions of Theorem 2 (2). We abuse the notation. Suppose also that the following four hold.
@@
    The critical set of $c_{1}$ is unbounded above, the restriction of $c_{1}$ to some interval of the form $\{t\geq b_{{\rm M},0}\}\subset\mathbb{R}$ is strongly decreasing.
@@
By the following exposition, our proof is completed.
The assumptions (14) are for us to have paths like the path in the proof of Proposition 1 departing from $p_{D}$ obtained by joining the set $\{(t,p_{D,2})\mid p_{D,1}\leq t\leq c_{2}(p_{D,2})\}$, which is a one-point set or a segment, and the path in $\overline{{D}_{m,c_{1},c_{2}}}-{D}_{m,c_{1},c_{2}}$ such that the restriction of ${\pi}_{2,1}$ there is injective and that the point $(p_{s,c_{2}},{(0)}_{j=1}^{m-1})\in{\mathbb{R}}^{2}\times{\mathbb{R}}^{m-1}={\mathbb{R}}^{m+1}$ is a critical point of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$.
@@
Suppose that (1) holds. In this case, the critical set of $c_{2}$ is bounded below and the restriction of $c_{2}$ to each interval of the form $\{t\leq b_{\rm m}\}\subset\mathbb{R}$ is not strongly increasing. We can see that the restriction of $c_{2}$ to some interval $\{t\leq b_{{\rm m},0}\}\subset\mathbb{R}$ is strongly decreasing and that there exist no critical point of $c_{1}$ on $\{t\leq b_{{\rm m},0}\}\subset\mathbb{R}$. We can also see that for $i=1,2$, the restriction of $c_{i}$ to the interval $\{t\leq b_{{\rm m},0}\}\subset\mathbb{R}$ is strongly decreasing.
The set $\{(x_{1},x_{2})\mid x_{2}\leq b_{{\rm m},0},c_{2}(b_{{\rm m},0})\leq x_{1}\leq c_{2}(x_{2}),c_{1}(x_{2})\leq x_{1}\leq c_{2}(x_{2})\}$ is connected and in $\overline{D_{c_{1},c_{2}}}$, and contains no critical point of $c_{i}$ for $i=1,2$. From this local property, we cannot construct paths as in the proofs of Propositions 1 and 2 and we can see that the Reeb space of the function ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ is regarded to be the Reeb graph and that it is not isomorphic to any finite or infinite graph. By a kind of symmetry, we can discuss similarly in the remaining three cases (24).

### v4

@@ Abstract.
Smooth, real analytic, or real algebraic (real polynomial) functions and maps. Reeb spaces. Graphs. Digraphs. Reeb (di)graphs.
@@
These tools or objects have been well-known as fundamental and important tools or objects in theory of Morse functions and more general differentiable functions with nice singularity ([15]). $R_{c}$ is, for a function $c$ which is not so wild mathematically, has the structure of a graph. For the Morse function case and Morse-Bott function case (on compact manifolds), see [3, 12] and for more general cases, see [1, 16, 17], for example. Especially, [1, Theorem 7.5], [16, Theorem 3.1], and [17, Theorems 2.1 and 2.8 and ”2 Reeb space and its graph structure”] are important as theory on topologies and graph structures of these spaces.
Hereafter, let ${\mathbb{R}}^{k}$ denote the $k$-dimensional Euclidean (real affine) space ($\mathbb{R}:={\mathbb{R}}^{1}$). Let ${\pi}_{k_{1},k_{2}}:{\mathbb{R}}^{k_{1}}\rightarrow{\mathbb{R}}^{k_{2}}$ denote the canonical projection mapping $x=(x_{1},x_{2})\in{\mathbb{R}}^{k_{1}-k_{2}}\times{\mathbb{R}}^{k_{2}}$ to $x_{1}$ with $k_{1}>k_{2}\geq 1$. A graph means a $1$-dimensional connected, locally finite, and locally compact CW complex such that the closure of each $1$-cell (edge) is homeomorphic to the $1$-dimensional disk $D^{1}:=\{x\in\mathbb{R}\mid x^{2}\leq 1\}$: 0-cell is a vertex. A graph with ends (with related notions) is defined by weakening the condition on edges as follows: the closure of each $1$-cell (edge) is homeomorphic to $D^{1}$ or a ray $\{x\geq 0\mid x\in\mathbb{R}\}$. For CW complexes, see [2] for example.
In the case of the smooth manifold $X$ with no boundary, $R_{c}$ is, in a certain nice case, a graph (with ends) whose vertex set consists of points for all critical contours of $c$. This is the so-called Reeb graph of $c$ and in most studies including presented studies above, this definition is adopted. We can see the Reeb graph as a digraph (with ends) by giving an orientation on each edge by the following rule: an edge $e$ is an oriented edge departing from (resp. entering) $v_{e}$ if the restriction of $\bar{c}$ to the closure of $e$ in the graph has the minimum (resp, maximum) at $v_{e}$.
@@
Recently the author has started to consider infinite graphs (with ends) as Reeb spaces and obtained several explicit cases. They are also constructed as, fully or densely, real algebraic or real analytic and based on the paper [6] and the preprints [8, 9]. Related preprints are [10, 11]. We ignore related non-trivial knowledge and discuss related arguments in self-contained ways. The present paper is also a kind of our new remark on examples of [10, 11], especially [10, Theorem 1]. There real algebraic or real analytic functions or maps are presented and answers to Problem 1 are presented as the restrictions of the canonical projections ${\pi}_{k_{1},k_{2}}$ to the zero sets of these nice smooth maps between the real affine spaces. We respect these examples and present answers from a kind of essentially generalized viewpoints.
In the next section, we prove our new result. This is for a generalization of the case presented in [10, Theorem 1 and problem 6] and in Example 1, this is explained again. Explicitly, we consider a natural smooth map onto the region surrounded by the graphs of two smooth real-valued functions $c_{1}$ and $c_{2}$ diverging to a same value at each infinity in ${\mathbb{R}}^{2}$, obtained via Theorem 1, and its composition with the canonical projection ${\pi}_{2,1}$. We investigate the Reeb space of this function.
@@
A proper map $c:X\rightarrow Y$ between topological spaces $X$ and $Y$ is a map whose preimage $c^{-1}(K)$ is compact for any each subset $K\subset X$. A non-proper map is one which may not be proper.
@@
This means that the ranks of the differential of the function are always $1$ on the zero set. The zero set is an $m$-dimensional smooth submanifold of ${\mathbb{R}}^{m+1}$ with no boundary. We can prove the remaining easily by the situation. ∎
@@
    Suppose that the intersection $\{\mathbb{R}-(\{q_{1}\}\bigcup\{q_{2}\})\}\bigcap(c_{1}(S(c_{1}))\bigcup c_{2}(S(c_{2})))$ is a discrete and closed subset of $\{\mathbb{R}-(\{q_{1}\}\bigcup\{q_{2}\})\}$ and that at least one function of $c_{1}$ and $c_{2}$ has at least one critical point. Suppose also the following.
@@
      As $t\in\mathbb{R}$ diverges to -$\infty$, then both the values $c_{1}(t)$ and $c_{2}(t)$ converge to $q_{1}\in\mathbb{R}$.
@@
      As $t\in\mathbb{R}$ diverges to $\infty$, then both the values $c_{1}(t)$ and $c_{2}(t)$ converge to $q_{2}\in\mathbb{R}$.
@@
      For $i=1,2$, each connected component $K_{c_{1},c_{2},q_{i},j}$ of the set ${{\pi}_{m+1,1}}^{-1}(q_{i})\bigcap X_{m,c_{1},c_{2}}$ has a small compact and connected neighborhood $N_{c_{1},c_{2},q_{i},j}$ in $X_{m,c_{1},c_{2}}$ consisting of contours of the the restriction ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ and exactly one contour in a level set of the form ${{\pi}_{m+1,1}}^{-1}(q_{1})\bigcap X_{m,c_{1},c_{2}}$ or ${{\pi}_{m+1,1}}^{-1}(q_{2})\bigcap X_{m,c_{1},c_{2}}$. The complementary set of $K_{c_{1},c_{2},q_{i},j}$ in the connected component $N_{c_{1},c_{2},q_{i},j}$ consists of finitely many connected sets and contains no critical contour of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$.
@@
We prove the statement (2). First, the critical sets of the functions $c_{1}$ and $c_{2}$ are discrete by the assumption. From the conditions (2a, 2b) and the structures of the maps and the manifolds, the restriction of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ to the preimage of $\mathbb{R}-(\{q_{1}\}\bigcup\{q_{2}\})$ is proper. The intersection $\{\mathbb{R}-(\{q_{1}\}\bigcup\{q_{2}\})\}\bigcap(c_{1}(S(c_{1}))\bigcup c_{2}(S(c_{2})))$ is a discrete and closed subset of $\{\mathbb{R}-(\{q_{1}\}\bigcup\{q_{2}\})\}$. By virtue of [16, 17], its Reeb space is regarded to be homeomorphic to a $1$-dimensional manifold or a graph which is not homeomorphic to a manifold. In addition, if at least one critical point of the function exists, then the Reeb space of the restriction of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ to the preimage of $\mathbb{R}-(\{q_{1}\}\bigcup\{q_{2}\})$ is regarded to be the Reeb graph. For [16, 17], related to this argument, mainly, consult [16, Theorem 3.1] and [17, Theorems 2.1 and 2.8 and ”2 Reeb space and its graph structure”]. Related to this theory, see also [1]. We consider the condition (2c). According to this, around each point in ${\bar{{\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}}}^{-1}(\{q_{1}\}\bigcup\{q_{2}\})$, the Reeb space is homeomorphic to either a $1$-dimensional manifold or a finite graph which is not homeomorphic to a manifold. This is also a Hausdorff space by the assumption that $N_{c_{1},c_{2},q_{i},j}$ in $X_{m,c_{1},c_{2}}$ contains exactly one contour of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ in level sets of the form ${{\pi}_{m+1,1}}^{-1}(q_{1})\bigcap X_{m,c_{1},c_{2}}$ or ${{\pi}_{m+1,1}}^{-1}(q_{2})\bigcap X_{m,c_{1},c_{2}}$. The Reeb space is also regarded to be the Reeb graph (if there exists a critical point of the function ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ there), locally. This also comes from the general theory above. We can easily see that two distinct contours of the function ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ are always separated by suitably chosen neighborhoods in $X_{m,c_{1},c_{2}}$ and the Reeb space of the function ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ is a Hausdorff space. It is also a locally connected space of course.
@@
Proposition 1.
In Theorem 1, if $S(c_{1})$ and $S(c_{2})$ are discrete and conditions of (2a, 2b) of Theorem 2 are satisfied, then the intersection $\{\mathbb{R}-(\{q_{1}\}\bigcup\{q_{2}\})\}\bigcap(c_{1}(S(c_{1}))\bigcup c_{2}(S(c_{2})))$ is a discrete and closed subset of $\mathbb{R}-(\{q_{1}\}\bigcup\{q_{2}\})$.
Proof.
A compact subset $K_{c_{1},c_{2}}$ in $\mathbb{R}-(\{q_{1}\}\bigcup\{q_{2}\})$ is also bounded and closed in $\mathbb{R}-(\{q_{1}\}\bigcup\{q_{2}\})$. In the preimage ${c_{1}}^{-1}({K_{c_{1},c_{2}}})\bigcup{c_{2}}^{-1}({K_{c_{1},c_{2}}})$, by the discreteness of the critical set of $c_{i}$ and the boundedness and compactness of ${c_{1}}^{-1}({K_{c_{1},c_{2}}})\bigcup{c_{2}}^{-1}({K_{c_{1},c_{2}}})$, $(S(c_{1})\bigcup S(c_{2}))\bigcap({c_{1}}^{-1}({K_{c_{1},c_{2}}})\bigcup{c_{2}}^{-1}({K_{c_{1},c_{2}}}))$ is finite and $(c_{1}(S(c_{1}))\bigcup c_{2}(S(c_{2})))\bigcap K_{c_{1},c_{2}}$ is discrete in $K_{c_{1},c_{2}}$. For each point $p$ of $\mathbb{R}-(\{q_{1}\}\bigcup\{q_{2}\})$, we have its compact neighborhood $K_{p,c_{1},c_{2}}\ni p$ in $\mathbb{R}-(\{q_{1}\}\bigcup\{q_{2}\})$. From this, we can see that $(\mathbb{R}-(\{q_{1}\}\bigcup\{q_{2}\}))\bigcap(c_{1}(S(c_{1}))\bigcup c_{2}(S(c_{2})))$ is a discrete and closed subset in $(\mathbb{R}-(\{q_{1}\}\bigcup\{q_{2}\}))\bigcap(c_{1}(S(c_{1}))\bigcup c_{2}(S(c_{2})))$. ∎
For such functions $c_{1}$ and $c_{2}$ of Proposition 1, Morse functions and real analytic functions are important.
@@
  • There exists an interval $\{t\mid a_{0}\leq t\leq b_{0}\}$ and $c$ has the maximum $c_{\rm M}$ there. Furthermore, there exist a sequence $\{a_{j}\}_{j=1}^{\infty}$ such that $a_{j}<a_{j+1}$ for each $j\geq 0$ and that as $j$ diverges to $\infty$, $a_{j}$ diverges to $\infty$ and another sequence $\{b_{j}\}_{j=1}^{\infty}$ such that $b_{j+1}<b_{j}$ for each $j\geq 0$ and that as $j$ diverges to $\infty$, $b_{j}$ diverges to -$\infty$ enjoying the following.
@@
    The case $c_{1}(x):=\frac{1}{2(x^{2}+1)}{\sin}^{2}x$ and $c_{2}(x):=\frac{1}{x^{2}+1}$ is an explicit case of Proposition 2 with the critical set of $c_{1}$ being unbounded, $c_{2}$ being (almost) unimodal, and $(q_{1},q_{2})=(0,0)$. This is also from [10, Theorem 1].
@@
    The case $c_{1}(x):=-\frac{1}{2(x^{2}+1)}{\sin}^{2}x$ and $c_{2}(x):=\frac{1}{x^{2}+1}$ seems to be an explicit case of Proposition 2 with the critical set of $c_{1}$ being unbounded, $c_{2}$ being almost unimodal, and $(q_{1},q_{2})=(0,0)$. However, the Reeb space is not homeomorphic to any graph with ends. Indeed, the condition (2c) from Theorem 2 is dropped and the set ${{\pi}_{2,1}}^{-1}(q_{i})\bigcap\overline{D_{c_{1},c_{2}}}$ and ${{\pi}_{m+1,1}}^{-1}(q_{i})\bigcap X_{m,c_{1},c_{2}}$ are connected and non-compact. the value of the quotient map $q_{{\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}}:X_{m,c_{1},c_{2}}\rightarrow R_{{\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}}$ at points in the set ${{\pi}_{m+1,1}}^{-1}(q_{i})\bigcap X_{m,c_{1},c_{2}}$, which is an unique element, is incident to infinitely many copies of $1$-cells, in the Reeb space. The Reeb space has the structure of a $1$-dimensional CW complex which is not locally finite or locally compact.
Hereafter, a real-valued smooth function $c_{I,\mathbb{R}}:I\rightarrow\mathbb{R}$ on $I\subset\mathbb{R}$, is strictly increasing (decreasing) if for any pair $(t_{1},t_{2})$ of real numbers $t_{1}<t_{2}$ in $I$, $c_{I,\mathbb{R}}(t_{1})<c_{I,\mathbb{R}}(t_{2})$ (resp. $c_{I,\mathbb{R}}(t_{1})>c_{I,\mathbb{R}}(t_{2})$) always holds.
In Proposition 2, if the critical set of $c_{2}$ is bounded below and $c_{2}$ is almost-unimodal, then the restriction of $c_{2}$ to some interval $\{t\leq b_{\rm m}\}\subset\mathbb{R}$ is strongly increasing, and if the critical set of $c_{2}$ is bounded above and $c_{2}$ is almost-unimodal, then the restriction of $c_{2}$ to some interval $\{t\geq b_{\rm M}\}\subset\mathbb{R}$ is strongly decreasing. There if the critical set of $c_{1}$ is bounded below and $c_{1}$ is almost-unimodal, then the restriction of $c_{1}$ to some interval $\{t\leq b_{\rm m}\}\subset\mathbb{R}$ is strongly increasing (for distinct numbers $t_{1}<t_{2}$ in the interval $c_{1}(t_{1})<-c_{1}(t_{2})$), and if the critical set of $c_{1}$ is bounded above and $c_{1}$ is almost-unimodal, then the restriction of $c_{1}$ to some interval $\{t\geq b_{\rm M}\}\subset\mathbb{R}$ is strongly decreasing (for distinct numbers $t_{1}<t_{2}$ in the interval $c_{1}(t_{1})>-c_{1}(t_{2})$).
@@
Let $c_{1}$ and $c_{2}$ be functions of Theorem 2 (2). We abuse the notation. Suppose also that all of the following four hold.
@@
    The critical set of $c_{1}$ is unbounded above, or the restriction of $c_{1}$ to some interval of the form $\{t\geq b_{{\rm M},0}\}\subset\mathbb{R}$ is strongly decreasing.
@@
By the following exposition, our proof is essentially completed.
The assumptions (14) enable us to find paths like the path in the proof of Proposition 2 departing from $p_{D}$ obtained by joining the set $\{(t,p_{D,2})\mid p_{D,1}\leq t\leq c_{2}(p_{D,2})\}$ ($\{(t,p_{D,2})\mid c_{2}(p_{D,2})\leq t\leq p_{D,1}\}$), which is a one-point set or a segment, and the path in $\overline{{D}_{m,c_{1},c_{2}}}-{D}_{m,c_{1},c_{2}}$ such that the restriction of ${\pi}_{2,1}$ to the resulting path is injective and that the point $(p_{s,c_{2}},{(0)}_{j=1}^{m-1})\in{\mathbb{R}}^{2}\times{\mathbb{R}}^{m-1}={\mathbb{R}}^{m+1}$ is a critical point of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$.
@@
Suppose that (1) holds.
In this case, the critical set of $c_{2}$ is bounded below and the restriction of $c_{2}$ to each interval of the form $\{t\leq b_{\rm m}\}\subset\mathbb{R}$ is not strongly increasing. We can see that the restriction of $c_{2}$ to some interval $\{t\leq b_{{\rm m},0}\}\subset\mathbb{R}$ is strongly decreasing and that there exist no critical point of $c_{1}$ on $\{t\leq b_{{\rm m},0}\}\subset\mathbb{R}$. We can also see that for $i=1,2$, the restriction of $c_{i}$ to the interval $\{t\leq b_{{\rm m},0}\}\subset\mathbb{R}$ is strongly decreasing.
The set $\{(x_{1},x_{2})\mid x_{2}\leq b_{{\rm m},0},c_{2}(b_{{\rm m},0})\leq x_{1}\leq c_{2}(x_{2}),c_{1}(x_{2})\leq x_{1}\leq c_{2}(x_{2})\}$ is connected and in $\overline{D_{c_{1},c_{2}}}$, and contains no critical point of $c_{i}$ for $i=1,2$. From this local property, we cannot construct paths as in the proofs of Propositions 2 and 3 and we can see that the Reeb space of the function ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ is regarded to be the Reeb graph and that it is not isomorphic to any finite or infinite graph.
By a kind of symmetry, we can discuss similarly in the remaining three cases (24).